# Einstein Fields: A Neural Perspective to Computational General Relativity

### v1

@@ A.3.5 Curvature tensors and scalars
[…]
@@
Another important tensor field of rank$(0,4)$ is the Weyl tensor, which is obtained as the “trace-free” part of the Riemann tensor. Physically, the Weyl tensor describes the tidal force experienced by a body when moving along geodesics, and quantifies the shape distortion a body experiences due to tidal forces (e.g., water tides caused by the gravitational pull of the moon). In an $n$-dimensional manifold it is defined as:
@@ A.3.6 Stress-energy-momentum tensor
The stress-energy-momentum tensor features on the right hand side of the Einstein field equations, and influences spacetime geometry by causing distortions on it.
@@ A.3.7 Einstein field equations
The Einstein field equations (EFEs) are a set of second-order non-linear PDEs containing geometry on the left hand side and the source on the right hand side. EFEs are obtained by combining all the differential geometric quantities from Equations ((51), (76), (77), (80)), together with the conservation laws for matter distribution of Equation (81), resulting in
@@
Here, $G_{\alpha\beta}:=R_{\alpha\beta}-\frac{1}{2}g_{\alpha\beta}R+\Lambda g_{\alpha\beta}$ is called the Einstein tensor and also satisfies the divergence-free condition $\nabla_{\alpha}G^{\alpha\beta}=0$, which, is a consequence of the IInd Bianchi identity of Equation (74b).
@@ Appendix B Exact Solutions of the Einstein Field Equations
This Appendix contains a detailed description of the exact solutions of EFEs corresponding to a class of metrics $g_{\mu\nu}$ that are solutions of Equation (82). While there exist several geometries that satisfy the EFEs, we shall consider three prominent geometries: the Schwarzschild metric, the Kerr metric, and gravitational waves. These solutions not only have a high theoretical and historical relevance, but are also of great interest in computational black hole astrophysics and gravitational wave and multi-messenger astronomy. From here on, we work in naturalized units by setting $c=G=1$.
@@ B.1.1 Coordinate systems for Schwarzschild metrics
$g_{\mu\nu}^{\text{Sph}}=\begin{pmatrix}-\bigg{(}1-\frac{2M}{r}\bigg{)}&0&0&0\\ 0&\bigg{(}1-\frac{2M}{r}\bigg{)}^{-1}&0&0\\ 0&0&r^{2}&0\\ 0&0&0&r^{2}\text{sin}^{2}\theta\end{pmatrix}\ .$ (85)
@@
where $\bar{g}_{\alpha\beta}$ is some background metric, $l_{\alpha}$ are the components of a null vector $\bm{\ell}$ with respect to the background metric and $V(x)$ is a scalar.
@@
Unlike the spherical coordinate form in Equation (84), $r=2M$ is not singular, hence removing the coordinate singularities. The metric tensor components read:
$g^{\text{KS}}_{\mu\nu}=\begin{pmatrix}-1+2M/r&\dfrac{2Mx}{r^{2}}&\dfrac{2My}{r^{2}}&\dfrac{2Mz}{r^{2}}\\[10.0pt] \dfrac{2Mx}{r^{2}}&1+\dfrac{2Mx^{2}}{r^{3}}&\dfrac{2Mxy}{r^{3}}&\dfrac{2Mxz}{r^{3}}\\[10.0pt] \dfrac{2My}{r^{2}}&\dfrac{2Mxy}{r^{3}}&1+\dfrac{2My^{2}}{r^{3}}&\dfrac{2Myz}{r^{3}}\\[10.0pt] \dfrac{2Mz}{r^{2}}&\dfrac{2Mxz}{r^{3}}&\dfrac{2Myz}{r^{3}}&1+\dfrac{2Mz^{2}}{r^{3}}\end{pmatrix}\ .$ (89)
@@ B.1.1.3 Ingoing (advanced) Eddington-Finkelstein coordinates
$g^{\text{EF}}_{\mu\nu}=\begin{pmatrix}-\left(1-\dfrac{2M}{r}\right)&1&0&0\\[10.0pt] 1&0&0&0\\[10.0pt] 0&0&r^{2}&0\\[10.0pt] 0&0&0&r^{2}\sin^{2}\vartheta\end{pmatrix}\ .$ (92)
@@ B.2.1 Coordinate systems for Kerr metric
$x$ $=\sqrt{r^{2}+a^{2}}\ \text{sin}\vartheta\ \text{cos}\phi$ (93a)
$y$ $=\sqrt{r^{2}+a^{2}}\ \text{sin}\vartheta\ \text{sin}\phi$ (93b)
@@
where the length scales are $a=\frac{J}{M}$ (angular momentum per unit mass), $\Sigma\equiv r^{2}+a^{2}\text{cos}^{2}\vartheta$, and $\Delta\equiv r^{2}-2Mr+a^{2}$. The Kerr curvature singularity occurs at $\Sigma:=r^{2}+a^{2}\text{cos}^{2}\vartheta=0$, implying $r=0$ and $\vartheta=\frac{\pi}{2}$. The metric tensor of Equation (94) is:
$g^{\text{BL}}_{\mu\nu}=\begin{pmatrix}-\left(1-\dfrac{2Mr}{\Sigma}\right)&0&0&-\dfrac{2Mar\sin^{2}\vartheta}{\Sigma}\\[10.0pt] 0&\dfrac{\Sigma}{\Delta}&0&0\\[10.0pt] 0&0&\Sigma&0\\[10.0pt] -\dfrac{2Mar\sin^{2}\vartheta}{\Sigma}&0&0&\left(r^{2}+a^{2}+\dfrac{2Ma^{2}r\sin^{2}\vartheta}{\Sigma}\right)\sin^{2}\vartheta\end{pmatrix}\ .$ (95)
@@
The Cartesian KS form of the Kerr metric is obtained by setting in Equation (87) $\bm{\ell}=\Big{(}1,\frac{rx+ay}{r^{2}+a^{2}},\frac{ry-ax}{r^{2}+a^{2}},\frac{z}{r}\Big{)}$ and $V=\frac{mr^{3}}{r^{4}+a^{2}z^{2}}$ in Equation  (87). The Kerr metric in Kerr coordinates are often used to write initial data for hydro simulations. The line-element in the Cartesian Kerr-Schild form reads (Teukolsky, 2015):
@@
[…]
@@
$+2(dv+a\ \text{sin}^{2}\theta d\tilde{\phi})(dr+a\ \text{sin}^{2}\theta d\tilde{\phi})$
@@
where the ingoing EF coordinates are related to the Boyer-Lindquist coordinates Equation (93) by the following transformation:
[…]
@@
$g^{\text{EF}}_{\mu\nu}=\begin{pmatrix}-\left(1-\dfrac{2Mr}{\Sigma}\right)&1&0&\dfrac{2Mar\sin^{2}\theta}{\Sigma}\\[10.0pt] 1&0&0&a\sin^{2}\theta\\[10.0pt] 0&0&\Sigma&0\\[10.0pt] \dfrac{2Mar\sin^{2}\theta}{\Sigma}&a\sin^{2}\theta&0&\left(r^{2}+a^{2}+\dfrac{2Ma^{2}r\sin^{2}\theta}{\Sigma}\right)\sin^{2}\theta\end{pmatrix}\ .$ (98)
@@
where, $r^{\text{ergo}}_{+}$ is the outer ergosphere, while $r^{\text{ergo}}_{-}$ demarcates the inner ergosphere. In Figure (13), the following regions of the Kerr metric are demarcated:
@@
B.2.2 Kretschmann scalar associated with Kerr solution
The nontrivial part of the Kretschmann invariant for the Kerr metric reads:
$r=0\ ,\text{and},\ \vartheta=\dfrac{\pi}{2}\ .$
@@ B.3 Gravitational waves
$\Box\overline{h}^{(\epsilon)}_{\alpha\beta}=-16\pi T_{\alpha\beta}\ .$
Here, $\overline{h}^{(\epsilon)}_{\alpha\beta}=h^{(\epsilon)}_{\alpha\beta}-\frac{1}{2}h^{(\epsilon)}\eta_{\alpha\beta}$ and $\Box\equiv\eta^{\alpha\beta}\partial_{\alpha}\partial_{\beta}$ is the d’Alembert operator (wave operator). It can be shown that the PDEs in Equation (B.3) produce gravitational wave solutions.
@@
$h_{+}$ and $h_{\times}$ are the amplitudes of the “$+$” (plus) polarization and “$\times$” (cross) polarization.
@@
$g_{\alpha\beta}=\eta_{\alpha\beta}+h^{\text{TT}}_{\alpha\beta}=\begin{pmatrix}-1&0&0&0\\[10.0pt] 0&1+h_{+}\cos\big{(}\omega(t-z)\big{)}&h_{\times}\cos\big{(}\omega(t-z)\big{)}&0\\[10.0pt] 0&h_{\times}\cos\big{(}\omega(t-z)\big{)}&1+h_{+}\cos\big{(}\omega(t-z)\big{)}&0\\[10.0pt] 0&0&0&1\end{pmatrix}\ .$ (101)
@@
[…]
where, $d\Omega=\sin\theta\,d\theta\,d\phi$ is the spherical volume element and ${}_{-2}\bar{Y}_{\ell m}$ denotes the complex conjugate of the $s=2-n=-2$ where ($n=4$ for $\Psi_{4}$) spin-weighted spherical harmonics. One typically carries out the integral over the 2-sphere $\mathbb{S}^{2}$ numerically on a finite angular grid.
@@
where the parameters $\ell,m$ are the familiar Laplace spherical harmonics (orbital-angular momentum and azimuthal indices), while $s$ is the additional spin-weight introduced by some underlying gauge group such as $U(1)$. We especially plot for the integers $s=-2,l=m=2$, since they are relevant for GWs and are depicted in Figure 14.
Figure 14: Spin weighted spherical harmonics for $s=-2$, and $l=2$ for $|m|\leq l$. The dominant contributions for the Weyl scalar $\Psi_{4}$ and the associated metric coefficients in the spherical harmonic basis $h^{2,\pm 2}(t)$ are shown.
@@ B.4.1 Coordinate systems for Minkowski metric
Setting $M\rightarrow 0$ in the Boyer-Lindquist form of the Kerr metric (94), the corresponding line element reduces to an unfamiliar “oblate-spheroidal" represtation:
@@
$\eta^{\text{BL}}_{\mu\nu}=\begin{pmatrix}-1&0&0&0\\ 0&\frac{r^{2}+a^{2}\text{cos}^{2}\vartheta}{r^{2}+a^{2}}&0&0\\ 0&0&r^{2}+a^{2}\text{cos}^{2}\vartheta&0\\ 0&0&0&(r^{2}+a^{2})\ \text{sin}^{2}\vartheta\end{pmatrix}\ .$ (109)
@@
    $\eta^{\text{EF}}_{\mu\nu}=\begin{pmatrix}-1&1&0&0\\ 1&0&0&0\\ 0&0&r^{2}&0\\ 0&0&0&r^{2}\ \text{sin}^{2}\theta\end{pmatrix}\ .$ (110)
@@
    $\eta^{\text{EF}}_{\mu\nu}=\begin{pmatrix}-1&1&0&0\\ 1&0&0&a\text{sin}^{2}\vartheta\\ 0&0&\Sigma&0\\ 0&a\text{sin}^{2}\vartheta&0&(r^{2}+a^{2})\ \text{sin}^{2}\vartheta\end{pmatrix}\ .$ (112)
@@ B.5 Training on non-trivial metric fields (distortions)
  • Spherical coordinates: obtained by subtracting (107) from (85):
    $g^{\text{Sph}}_{\mu\nu}-\eta^{\text{Sph}}_{\mu\nu}=\begin{pmatrix}\dfrac{r_{s}}{r}&0&0&0\\[10.0pt] 0&\dfrac{r_{s}}{(r-r_{s})}&0&0\\[10.0pt] 0&0&0&0\\[10.0pt] 0&0&0&0\ \end{pmatrix}\ .$ (113)
  • Kerr-Schild coordinates obtained by subtracting $\eta_{\mu\nu}=\text{diag}(-1,+1,+1,+1)$ from (89):
    $g^{\text{KS}}_{\mu\nu}-\eta_{\mu\nu}=\begin{pmatrix}\dfrac{2M}{r}&\dfrac{2Mx}{r^{2}}&\dfrac{2My}{r^{2}}&\dfrac{2Mz}{r^{2}}\\[10.0pt] \dfrac{2Mx}{r^{2}}&\dfrac{2Mx^{2}}{r^{3}}&\dfrac{2Mxy}{r^{3}}&\dfrac{2Mxz}{r^{3}}\\[10.0pt] \dfrac{2My}{r^{2}}&\dfrac{2Mxy}{r^{3}}&\dfrac{2My^{2}}{r^{3}}&\dfrac{2Myz}{r^{3}}\\[10.0pt] \dfrac{2Mz}{r^{2}}&\dfrac{2Mxz}{r^{3}}&\dfrac{2Myz}{r^{3}}&\dfrac{2Mz^{2}}{r^{3}}\end{pmatrix}\ .$ (114)
  • Ingoing Eddington-Finkelstein coordinates obtained by subtracting (110) from (92):
    $g^{\text{EF}}_{\mu\nu}-\eta^{\text{EF}}_{\mu\nu}=\begin{pmatrix}\dfrac{r_{s}}{r}&0&0&0\\[10.0pt] 0&0&0&0\\[10.0pt] 0&0&0&0\\[10.0pt] 0&0&0&0\end{pmatrix}\ .$ (115)
@@
  • Boyer-Lindquist coordinates: obtained by subtracting (109) from (95):
    $g^{\text{BL}}_{\mu\nu}-\eta^{\text{BL}}_{\mu\nu}=\begin{pmatrix}\dfrac{2Mr}{\Sigma}&0&0&\dfrac{2Mar\sin^{2}\theta}{\Sigma}\\[10.0pt] 0&\dfrac{2M\Sigma}{r\Delta}&0&0\\[10.0pt] 0&0&0&0\\[10.0pt] \dfrac{2Mar\sin^{2}\theta}{\Sigma}&0&0&\dfrac{2Ma^{2}r\sin^{4}\vartheta}{\Sigma}\ \end{pmatrix}\ .$ (116)
  • Kerr-Schild coordinates: obtained by subtracting $\eta=\text{diag }(-1,+1,+1,+1)$ from Equation (96):
  • Eddington-Finkelstein coordinates: obtained by subtracting (112) from (98):
    $g^{\text{EF}}_{\mu\nu}-\eta^{\text{EF}}_{\mu\nu}=\begin{pmatrix}\dfrac{2Mr}{\Sigma}&0&0&\dfrac{2Mar\sin^{2}\theta}{\Sigma}\\[10.0pt] 0&0&0&0\\[10.0pt] 0&0&0&0\\[10.0pt] \dfrac{2Mar\sin^{2}\theta}{\Sigma}&0&0&\dfrac{2Ma^{2}r\sin^{4}\vartheta}{\Sigma}\ \end{pmatrix}\ .$ (117)

### v2

@@ A.3.5 Curvature tensors and scalars
$R_{\alpha\beta\gamma\delta}+R_{\alpha\gamma\delta\beta}+R_{\alpha\delta\beta\gamma}$ $=0\ (\text{Bianchi Identity I)}\ ,$ (59a)
$\nabla_{\alpha}R_{\beta\gamma\delta\sigma}+\nabla_{\beta}R_{\gamma\alpha\delta\sigma}+\nabla_{\gamma}R_{\alpha\beta\delta\sigma}$ $=0\ (\text{Bianchi Identity II)}\ .$ (59b)
@@
Another important tensor field of rank(0,4)(0,4) is the Weyl tensor, which is obtained as the “trace-free” part of the Riemann tensor. Physically, the Weyl tensor describes the tidal force experienced by a body when moving along geodesics, and quantifies the shape distortion a body experiences due to tidal forces (e.g., water tides caused by the gravitational pull of the moon). In an $n$-dimensional manifold it is defined as:
@@ A.3.6 Stress-energy-momentum tensor
The stress-energy-momentum tensor features on the right hand side of the Einstein field equations (EFEs), and influences spacetime geometry by causing distortions on it.
@@ A.3.7 Einstein field equations
The EFEs are a set of second-order non-linear PDEs containing geometry on the left hand side and the source on the right hand side. EFEs are obtained by combining all the differential geometric quantities from Eqs. (36, 61, 62, 65), together with the conservation laws for matter distribution of Eq. (66), resulting in
@@
Here, $G_{\alpha\beta}:=R_{\alpha\beta}-\frac{1}{2}g_{\alpha\beta}R+\Lambda g_{\alpha\beta}$ is called the Einstein tensor and also satisfies the divergence-free condition $\nabla_{\alpha}G^{\alpha\beta}=0$, which, is a consequence of the IInd Bianchi identity of Eq. (59b).
@@ Appendix B Exact Solutions of the Einstein Field Equations
This Appendix contains a detailed description of the exact solutions of EFEs corresponding to a class of metrics $g_{\mu\nu}$ that are solutions of Eq. (67). While there exist several geometries that satisfy the EFEs, we shall consider three prominent geometries: the Schwarzschild metric, the Kerr metric, and gravitational waves. These solutions not only have a high theoretical and historical relevance, but are also of great interest in computational black hole astrophysics and gravitational wave and multi-messenger astronomy. From here on, we work in naturalized units by setting $c=G=1$.
@@ B.1.1 Coordinate systems for Schwarzschild metrics
$g_{\mu\nu}^{\text{Sph}}=\begin{pmatrix}-\bigg(1-\frac{2M}{r}\bigg)&0&0&0\\ 0&\bigg(1-\frac{2M}{r}\bigg)^{-1}&0&0\\ 0&0&r^{2}&0\\ 0&0&0&r^{2}\text{sin}^{2}\theta\end{pmatrix}\$. (70)
@@
[…]
@@
Unlike the spherical coordinate form in Eq. (D.0.1.1), $r=2M$ is not singular, hence removing the coordinate singularities. The metric tensor components read:
$g^{\text{KS}}_{\mu\nu}=\begin{pmatrix}-1+2M/r&\dfrac{2Mx}{r^{2}}&\dfrac{2My}{r^{2}}&\dfrac{2Mz}{r^{2}}\\[10.0pt] \dfrac{2Mx}{r^{2}}&1+\dfrac{2Mx^{2}}{r^{3}}&\dfrac{2Mxy}{r^{3}}&\dfrac{2Mxz}{r^{3}}\\[10.0pt] \dfrac{2My}{r^{2}}&\dfrac{2Mxy}{r^{3}}&1+\dfrac{2My^{2}}{r^{3}}&\dfrac{2Myz}{r^{3}}\\[10.0pt] \dfrac{2Mz}{r^{2}}&\dfrac{2Mxz}{r^{3}}&\dfrac{2Myz}{r^{3}}&1+\dfrac{2Mz^{2}}{r^{3}}\end{pmatrix}\$. (74)
@@ B.1.1.3 Ingoing (advanced) Eddington-Finkelstein coordinates
$g^{\text{EF}}_{\mu\nu}=\begin{pmatrix}-\left(1-\dfrac{2M}{r}\right)&1&0&0\\[10.0pt] 1&0&0&0\\[10.0pt] 0&0&r^{2}&0\\[10.0pt] 0&0&0&r^{2}\sin^{2}\vartheta\end{pmatrix}\$. (77)
@@ B.2.1 Coordinate systems for Kerr metric
$x$ $=\sqrt{r^{2}+a^{2}}\ \text{sin}\vartheta\ \text{cos}\phi\$(78a)
$y$ $=\sqrt{r^{2}+a^{2}}\ \text{sin}\vartheta\ \text{sin}\phi\$(78b)
@@
where the length scales are $a=\frac{J}{M}$ (angular momentum per unit mass), $\Sigma\equiv r^{2}+a^{2}\text{cos}^{2}\vartheta$, and $\Delta\equiv r^{2}-2Mr+a^{2}$. The Kerr curvature singularity occurs at $\Sigma:=r^{2}+a^{2}\text{cos}^{2}\vartheta=0$, implying $r=0$ and $\vartheta=\frac{\pi}{2}$. The metric tensor of Eq. (79) is:
$g^{\text{BL}}_{\mu\nu}=\begin{pmatrix}-\left(1-\dfrac{2Mr}{\Sigma}\right)&0&0&-\dfrac{2Mar\sin^{2}\vartheta}{\Sigma}\\[10.0pt] 0&\dfrac{\Sigma}{\Delta}&0&0\\[10.0pt] 0&0&\Sigma&0\\[10.0pt] -\dfrac{2Mar\sin^{2}\vartheta}{\Sigma}&0&0&\left(r^{2}+a^{2}+\dfrac{2Ma^{2}r\sin^{2}\vartheta}{\Sigma}\right)\sin^{2}\vartheta\end{pmatrix}\$. (80)
@@
The Cartesian KS form of the Kerr metric is obtained by setting in Eq. (72) $\bm{\ell}=\Big(1,\frac{rx+ay}{r^{2}+a^{2}},\frac{ry-ax}{r^{2}+a^{2}},\frac{z}{r}\Big)$ and $V=\frac{mr^{3}}{r^{4}+a^{2}z^{2}}$ in Eq. (72). The Kerr metric in Kerr coordinates are often used to write initial data for hydro simulations. The line-element in the Cartesian Kerr-Schild form reads (Teukolsky, 2015):
@@
In the BL coordinates the ring singularity for Kerr exists at $r=0\ \&\ \vartheta=\frac{\pi}{2}$, translating to z=0(equatorial-plane)z=0\ ($\text${equatorial-plane}), and the ring occurring at $x^{2}+y^{2}=a^{2}$. In contrast, the KS representation is devoid of coordinate singularities, making it suitable to work in numerics, especially around the event-horizons.
@@
[…]
@@
where the ingoing EF coordinates are related to the Boyer-Lindquist coordinates Eq. (78) by the following transformation:
$v$ $=t+\int\dfrac{(r^{2}+a^{2})}{\Delta}dr\ ,$
@@
$g^{\text{EF}}_{\mu\nu}=\begin{pmatrix}-\left(1-\dfrac{2Mr}{\Sigma}\right)&1&0&\dfrac{2Mar\sin^{2}\theta}{\Sigma}\\[10.0pt] 1&0&0&a\sin^{2}\theta\\[10.0pt] 0&0&\Sigma&0\\[10.0pt] \dfrac{2Mar\sin^{2}\theta}{\Sigma}&a\sin^{2}\theta&0&\left(r^{2}+a^{2}+\dfrac{2Ma^{2}r\sin^{2}\theta}{\Sigma}\right)\sin^{2}\theta\end{pmatrix}\$. (83)
@@
where, $r^{\text{ergo}}_{+}$ is the outer ergosphere, while $r^{\text{ergo}}_{-}$ demarcates the inner ergosphere. In Figure 8, the following regions of the Kerr metric are demarcated:
@@ B.3 Gravitational waves
$\Box\overline{h}^{(\epsilon)}_{\alpha\beta}=-16\pi T_{\alpha\beta}\ .$ (84)
Here, $\overline{h}^{(\epsilon)}_{\alpha\beta}=h^{(\epsilon)}_{\alpha\beta}-\frac{1}{2}h^{(\epsilon)}\eta_{\alpha\beta}$ and $\Box\equiv\eta^{\alpha\beta}\partial_{\alpha}\partial_{\beta}$ is the d’Alembert operator (wave operator). It can be shown that the PDEs in Eq. (84) produce gravitational wave solutions.
@@
$h_{+}$ and $h_{\times}$ are the amplitudes of the “++” (plus) polarization and $\times$” (cross) polarization.
@@
$g_{\alpha\beta}=\eta_{\alpha\beta}+h^{\text{TT}}_{\alpha\beta}=\begin{pmatrix}-1&0&0&0\\[10.0pt] 0&1+h_{+}\cos\big(\omega(t-z)\big)&h_{\times}\cos\big(\omega(t-z)\big)&0\\[10.0pt] 0&h_{\times}\cos\big(\omega(t-z)\big)&1+h_{+}\cos\big(\omega(t-z)\big)&0\\[10.0pt] 0&0&0&1\end{pmatrix}\$. (86)
@@
$\dfrac{r}{M}h^{\ell m}(t)=\int_{0}^{2\pi}\int_{0}^{\pi}h(t,r,\theta,\phi)\,{}_{-2}\bar{Y}_{\ell m}(\theta,\phi)\,d\Omega$ (89)
where, $d\Omega=\sin\theta\,d\theta\,d\phi$ is the spherical volume element and $_{-2}\bar{Y}_{\ell m}$ denotes the complex conjugate of the s=2n=2s=2-n=-2 where ($n=4$ for $\Psi_{4}$) spin-weighted spherical harmonics. One typically carries out the integral over the 2-sphere $\mathbb{S}^{2}$ numerically on a finite angular grid.
@@
where the parameters $\ell,m$ are the familiar Laplace spherical harmonics (orbital-angular momentum and azimuthal indices), while $s$ is the additional spin-weight introduced by some underlying gauge group such as $U(1)$. We especially plot for the integers s=2,l=m=2s=-2,l=m=2, since they are relevant for GWs and are depicted in Figure 9.
Figure 9: Spin weighted spherical harmonics for s=2s=-2, and $l=2$ for $|m|\leq l$. The dominant contributions for the Weyl scalar $\Psi_{4}$ and the associated metric coefficients in the spherical harmonic basis $h^{2,\pm 2}(t)$ are shown.
@@ B.4.1 Coordinate systems for Minkowski metric
Setting $M\rightarrow 0$ in the Boyer-Lindquist form of the Kerr metric Eq. (79), the corresponding line element reduces to an unfamiliar “oblate-spheroidal" represtation:
@@
$\eta^{\text{BL}}_{\mu\nu}=\begin{pmatrix}-1&0&0&0\\ 0&\frac{r^{2}+a^{2}\text{cos}^{2}\vartheta}{r^{2}+a^{2}}&0&0\\ 0&0&r^{2}+a^{2}\text{cos}^{2}\vartheta&0\\ 0&0&0&(r^{2}+a^{2})\ \text{sin}^{2}\vartheta\end{pmatrix}\$. (94)
@@
    $\eta^{\text{EF}}_{\mu\nu}=\begin{pmatrix}-1&1&0&0\\ 1&0&0&0\\ 0&0&r^{2}&0\\ 0&0&0&r^{2}\ \text{sin}^{2}\theta\end{pmatrix}\$. (95)
@@
    $\eta^{\text{EF}}_{\mu\nu}=\begin{pmatrix}-1&1&0&0\\ 1&0&0&a\text{sin}^{2}\vartheta\\ 0&0&\Sigma&0\\ 0&a\text{sin}^{2}\vartheta&0&(r^{2}+a^{2})\ \text{sin}^{2}\vartheta\end{pmatrix}\$. (97)
@@ B.5 Training on non-trivial metric fields (distortions)
  • Spherical coordinates: obtained by subtracting Eq. (92) from Eq. (70):
    $g^{\text{Sph}}_{\mu\nu}-\eta^{\text{Sph}}_{\mu\nu}=\begin{pmatrix}\dfrac{r_{s}}{r}&0&0&0\\[10.0pt] 0&\dfrac{r_{s}}{(r-r_{s})}&0&0\\[10.0pt] 0&0&0&0\\[10.0pt] 0&0&0&0\ \end{pmatrix}\$. (98)
  • Kerr-Schild coordinates obtained by subtracting $\eta_{\mu\nu}=\text{diag}(-1,+1,+1,+1)$ from Eq. (74):
    $g^{\text{KS}}_{\mu\nu}-\eta_{\mu\nu}=\begin{pmatrix}\dfrac{2M}{r}&\dfrac{2Mx}{r^{2}}&\dfrac{2My}{r^{2}}&\dfrac{2Mz}{r^{2}}\\[10.0pt] \dfrac{2Mx}{r^{2}}&\dfrac{2Mx^{2}}{r^{3}}&\dfrac{2Mxy}{r^{3}}&\dfrac{2Mxz}{r^{3}}\\[10.0pt] \dfrac{2My}{r^{2}}&\dfrac{2Mxy}{r^{3}}&\dfrac{2My^{2}}{r^{3}}&\dfrac{2Myz}{r^{3}}\\[10.0pt] \dfrac{2Mz}{r^{2}}&\dfrac{2Mxz}{r^{3}}&\dfrac{2Myz}{r^{3}}&\dfrac{2Mz^{2}}{r^{3}}\end{pmatrix}\$. (99)
  • Ingoing Eddington-Finkelstein coordinates obtained by subtracting Eq. (95) from Eq. (77):
    $g^{\text{EF}}_{\mu\nu}-\eta^{\text{EF}}_{\mu\nu}=\begin{pmatrix}\dfrac{r_{s}}{r}&0&0&0\\[10.0pt] 0&0&0&0\\[10.0pt] 0&0&0&0\\[10.0pt] 0&0&0&0\end{pmatrix}\$. (100)
@@
  • Boyer-Lindquist coordinates: obtained by subtracting Eq. (94) from Eq. (80):
    $g^{\text{BL}}_{\mu\nu}-\eta^{\text{BL}}_{\mu\nu}=\begin{pmatrix}\dfrac{2Mr}{\Sigma}&0&0&\dfrac{2Mar\sin^{2}\theta}{\Sigma}\\[10.0pt] 0&\dfrac{2M\Sigma}{r\Delta}&0&0\\[10.0pt] 0&0&0&0\\[10.0pt] \dfrac{2Mar\sin^{2}\theta}{\Sigma}&0&0&\dfrac{2Ma^{2}r\sin^{4}\vartheta}{\Sigma}\ \end{pmatrix}\$. (101)
  • Kerr-Schild coordinates: obtained by subtracting $\eta=\text{diag }(-1,+1,+1,+1)$ from Eq. (81):
  • Eddington-Finkelstein coordinates: obtained by subtracting Eq. (97) from Eq. (83):
    $g^{\text{EF}}_{\mu\nu}-\eta^{\text{EF}}_{\mu\nu}=\begin{pmatrix}\dfrac{2Mr}{\Sigma}&0&0&\dfrac{2Mar\sin^{2}\theta}{\Sigma}\\[10.0pt] 0&0&0&0\\[10.0pt] 0&0&0&0\\[10.0pt] \dfrac{2Mar\sin^{2}\theta}{\Sigma}&0&0&\dfrac{2Ma^{2}r\sin^{4}\vartheta}{\Sigma}\ \end{pmatrix}\$. (102)